OpenStudy (anonymous):

Solve 6x2 + x - 1 = 0

OpenStudy (manish00333):

do you know how you would normally solve a quadratic equation?

OpenStudy (manish00333):

a+b=1 a x b = 6 x -1 = -6 a = 3 b=-2 6x^2 + 3x -2x - 1 = 0 3x(2x+1) -1(2x+1) (2x+1)(3x-1)=0 x=-1/2 x=1/3

OpenStudy (anonymous):

i got that but when u feill them in they dont work..

OpenStudy (manish00333):

they work! 6(1/3)^2 + 1/3 - 1 = 0 6(1/9) + 1/3 - 1 = 0 2/3 + 1/3 -1 = 0 1 - 1 = 0 0 = 0 6(-1/2)^2 - 1/2 - 1 = 0 6(1/4) - 1/2 -1 = 0 3/2 - 1/2 -1 = 0 1 - 1 = 0 0 = 0